# Why machine learning models fail to fully capture epistemic uncertainty

### v1

@@ 2.2 Uncertainty from a bias-variance perspective
$\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})=\underbrace{\mathbb{E}_{\mathcal{D}_{N}}\big{[}\text{Var}_{\gamma}(\hat{y}|\boldsymbol{x})|\mathcal{D}_{N}\big{]}}_{\text{procedural uncertainty}}+\underbrace{\text{Var}_{\mathcal{D}_{N}}(\mathbb{E}_{\gamma}[\hat{y}|\boldsymbol{x}]|\mathcal{D}_{N})}_{\text{data uncertainty}}.$ (4)
@@
The third term reflects the squared bias, which is the difference between the expected prediction $\hat{y}$ under data and model uncertainty and the true conditional expectation $\mathbb{E}_{y|\boldsymbol{x}}\big{[}y|\boldsymbol{x}\big{]}$. For universal approximators, such as many types of deep neural networks, the model uncertainty is zero [27, 28, 29]. Such models are asymptotically unbiased, but in the finite data regime, a bias is unavoidable to make learning possible [24]. For more restrictive hypothesis spaces, the bias even does not disappear when the training dataset becomes infinitely large. Thus, in both cases, the bias term cannot be ignored when analysing epistemic uncertainty.
@@ 3 Evaluating epistemic uncertainty via statistical simulations
Quantifying epistemic uncertainty objectively is challenging, and evaluating the correctness of such quantifications is equally difficult. While calibration tests exist for aleatoric uncertainty in probabilistic models [30], epistemic uncertainty lacks established evaluation protocols and is often assessed indirectly through downstream tasks like out-of-distribution detection [5, 15, 31], robustness to adversarial attacks [32, 33], and active learning [34]. We believe that correctness of second-order epistemic uncertainty representations can only be evaluated through statistical simulations, in which bias and variance can be analysed individually. To this end, Jürgens et al. [17] recently introduced the notion of a reference distribution, which is the theoretical distribution over first-order models if training data could be resampled. Here we further extend this notion by taking the variance due to procedural uncertainty into account:
Definition 3.1.
@@
[…]
@@
Intuitively, the reference distribution describes the uncertainty in the first-order parameters, induced by having a limited dataset and procedural uncertainty due to different weight initializations. In simulation studies, one can estimate the reference distribution by drawing $n_{d}$ training datasets $\mathcal{D}_{N_{1}},\dots,\mathcal{D}_{N_{n_{d}}}$ and $n_{\gamma}$ procedural parameters $\gamma_{1},\dots,\gamma_{n_{\gamma}}$ from $P^{N}$ and $P^{\gamma}$, respectively, and calculating the empirical distribution function.
The variance of the reference distribution describes the variance due to limited data expressivity, as well as the procedural variance, and can be decomposed by means of the law of total variance in Eqn. 4. In addition, in statistical simulations, where the data-generating process $p(y|\boldsymbol{x})$ is known, the bias of the reference distribution can also be measured.
Example 3.2 (An illustrative statistical simulation).
We illustrate this argument by adapting the synthetic regression setting from Rossellini et al. [35]. Precisely, we generate datasets $\mathcal{D}_{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\sim Beta(1.2,0.5)$, $y_{i}=\sin(\frac{1}{5(x_{i}+0.16)^{3}})+\epsilon_{i}$, where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(x_{i}))$ with $\sigma^{2}(x)=x^{4}$. The setup can be seen in Figure 1. This dataset highlights the differences between aleatoric and epistemic uncertainty: for low $x$ values, the complexity of the underlying data-generating function is high, while data is limited, leading to high epistemic uncertainty. On the other hand, for $x$ close to $1$, aleatoric uncertainty is high, while simultaneously having more data points and a less complex underlying function, implying low epistemic uncertainty. In order to estimate the reference distribution we train models with $n_{d}$ training datasets and $n_{\gamma}$ random weight initializations. As the underlying model class for obtaining the individual risk minimizers $\boldsymbol{\hat{\theta}}_{\gamma,N}$, we use an MLP. Details about the experimental setup are discussed in Appendix A.1111The implementation for all our examples can be found in this GitHub repository.. Individual estimates for data- and procedural uncertainty can be obtained by the decomposition formula in Eqn.4. The bias can be calculated as the absolute distance between the mean prediction and the underlying ground truth conditional expectation.
@@
Figure 3: Estimated procedural and data uncertainty of the reference distribution for increasing sample sizes. The reference distribution is estimated by resampling training data and procedural parameters, using $n_{\gamma}=10$ and $n_{d}=20$. Procedural and data uncertainty are calculated using the decomposition formula in Eqn. 4. Even when the model predictions (complete line) have an important deviation (bias) from the ground-truth expected value (dashed line), the epistemic uncertainty estimate decreases with more data.
@@
Figure 3 shows the obtained reference distribution for different training dataset sizes. First, we see that the total variance of the reference distribution decreases when sample size increases, supporting the claim that the reference distribution properly represents data uncertainty. Second, we see that the estimate of the procedural uncertainty also decreases with an increasing sample size, which is not unexpected, because initial procedural configurations become less important when more data is observed. Third, in the region with high epistemic uncertainty, the reference distribution has a very large bias, due to the underlying complexity of the data generating function. The bias is in fact so high that aleatoric uncertainty cannot be correctly estimated anymore This last finding confirms our central position: Any claim to properly represent epistemic uncertainty must examine not only how different sources of uncertainty are modeled, but also whether bias is systematically distorting the partitioning of uncertainty. When the bias is high, estimates of both epistemic and aleatoric uncertainty can be completely misleading.
@@ 4 Which part of epistemic uncertainty do current methods capture?
Departing from a Bayesian perspective, we first describe in this section how various second-order methods claim to obtain disentangled estimates for aleatoric and epistemic uncertainty. Secondly, we analyse two representative methods that approximate the posterior distribution over the parameters in more detail, illustrating that they tackle different sources of epistemic uncertainty. In the last part of this section, we present experimental results for eight second-order methods, highlighting that all of them have difficulties to model the different sources of epistemic uncertainty in a complete manner.
@@ 4.1 Second-order uncertainty quantification methods
Bayesian methods form principled approach to quantify and interpret epistemic uncertainty by modelling a prior and posterior distribution over the model parameters and computing the second-order distribution via Eqn. 2. For the regression setup that we consider in this paper, a common approach is to use Heteroscedastic Gaussian Processes (GP, [23, 36]), which place priors on the parameters $f(\boldsymbol{x})$ and $\sigma^{2}(\boldsymbol{x})$, thereby analytically computing the predictive distribution using two Gaussian processes, $f(\boldsymbol{x}),\log\sigma^{2}(\boldsymbol{x})\sim GPs$. For Bayesian neural networks, which have gained popularity in recent years, the posterior and second-order distribution cannot be analytically computed, due to the intractability of $p(\boldsymbol{\phi}|\mathcal{D}_{N})\propto p(\mathcal{D}_{N}|\boldsymbol{\phi})p(\boldsymbol{\phi})$ induced by the hypothesis space [1]. Due to the impossibility of performing exact inference in Bayesian neural networks, one typically obtains $d$ different samples $\{\boldsymbol{\phi}_{j}\}_{j=1}^{d}$ from $p(\boldsymbol{\phi}|\mathcal{D}_{N})$ by use of approximation techniques that aim to find a posterior distribution via sampling methods. Here, the second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ of Eqn. 2 is
is then approximated by Monte Carlo integration:
@@
In the experiments below, we employ three of the most common methods to obtain $\{\boldsymbol{\phi}_{i}\}_{j=1}^{M}$. Variational inference, (VI, [37, 38]) considers a variational (simpler) distribution $q(\boldsymbol{\phi})$ and optimizes the lower bound of the marginal likelihood (ELBO). Hamiltonian Monte Carlo (HMC) uses Hamiltonian dynamics to sample $\{\boldsymbol{\phi}_{j}\}$ from the true posterior. Laplace approximation (LA, [39]) builds a Taylor expansion around the maximum-a-posteriori estimate $\boldsymbol{\phi}_{\text{MAP}}$, yielding a Gaussian approximation.
Ensemble methods are also often interpreted as performing approximate Bayesian inference by finding multiple modes of $p(\boldsymbol{\phi}|\mathcal{D}_{N})$ after training multiple independent models. [40, 41]. In this work, we consider three types of ensembles. Deep Ensembles (DE, [9]), train $d$ networks on the same dataset $\mathcal{D}_{N}$, but with different random initialization of the weights $\boldsymbol{\phi}$. Similarly, in Bootstrap Ensembles (BE), $M$ networks are trained using different weight initializations, but each network uses a bootstrap sample $\mathcal{D}^{(j)}\subseteq\mathcal{D}_{N}$. While Bootstrap Ensembles have been shown to perform worse than training $d$ networks using Deep Ensembles [9], we include them in our analysis to estimate the effect of data uncertainty on the epistemic uncertainty estimates. Monte Carlo Dropout (MCD, [42]) can also be understood as a Bayesian approximation, where one trains a single neural network with dropout rate $\rho$ and performs multiple stochastic forward passes at inference time with the same dropout rate. In a different direction, evidential deep learning methods learn directly the second-order distribution over the predictive parameters $\theta$. Deep evidential regression (DEL, [12]) learns hyperparameters of a conjugate prior (e.g. Normal-Inverse-Gamma for the univariate normal distribution) directly using a single network trained by second-order loss minimization.
From a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$, one can decompose its representation of uncertainty into an estimate of aleatoric and epistemic uncertainty. Proposed measures focus on entropy [3], distance to reference sets [10], or proper scoring rules [43]. However, the most commonly used one for regression is arguably the variance-based decomposition [3, 10], which considers the variance of the prediction $\hat{y}$, obtained via model averaging as in Eqn. 2. The latter can be decomposed into (estimates of) an aleatoric and an epistemic part:
@@ 4.2 Two representative cases
Example 4.1 (Deep Ensembles purely capture procedural uncertainty).
Let $\gamma$ denote the random seed (weight initialization, data ordering etc.) that differentiates each ensemble member’s training, which in turn are trained on the same dataset $\mathcal{D}_{N}$. For a Deep Ensemble of size $d$, each ensemble member’s predictive parameters are indexed over the set of random seeds $\{\gamma_{1},\dots,\gamma_{d}\}$, obtaining a set of predictions $\{\boldsymbol{\theta}_{\gamma_{i}}(\boldsymbol{x})=(f_{\gamma_{i}}(\boldsymbol{x}),\sigma_{\gamma_{i}}^{2}(\boldsymbol{x}))\}_{i=1}^{d}$. The empirical second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ can be approximated as a mixture of equally weighted Dirac measures
@@
where $\hat{y}_{N,\gamma}(\boldsymbol{x}):=f_{\gamma}(\boldsymbol{x})$ corresponds to the prediction of an individual ensemble member trained with random initialization $\gamma$ on $\mathcal{D}_{N}$ and $\hat{y}_{N}(\boldsymbol{x}):=\hat{f}(\boldsymbol{x})$ of the aggregated ensembles. Hence, we see that the predicted epistemic uncertainty corresponds to the procedural uncertainty as illustrated in Figure 1 a). Model uncertainty through bias as well as data uncertainty are not represented.
@@
Example 4.2 (DER may fail to provide to faithful epistemic uncertainty estimates).

### v2

@@ 2.2 Uncertainty from a generalized bias-variance perspective
$+\underbrace{\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(\bar{p}^{*}(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x})\big]}_{\text{generalized variance}},$ (3)
where $\bar{p}^{*}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{p\in\mathbb{P}(\mathcal{Y})}D_{\ell}(p,\hat{p}_{N,\gamma}(y|\boldsymbol{x}))$ the dual mean of $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ under data- and procedural uncertainty. The first term in Eqn. 2.2 corresponds to the intrinsic variability within the data-generating process, and hence to aleatoric uncertainty. The last term reflects the generalized variance of the predictive distribution under data and procedural uncertainty, and ca,n under suitable conditions of the loss function $\ell$ [31] be further decomposed by means of the conditional bias-variance decomposition for Bregman divergences [30] as
$\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(\bar{p}^{*}(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x})\big]$ $=\underbrace{\mathbb{E}_{\mathcal{D}_{N}}\big[D_{\ell}(\bar{p}^{*}(y|\boldsymbol{x}),\bar{p}_{|\mathcal{D}_{N}}^{*}(y|\boldsymbol{x}))\big]}_{\text{procedural uncertainty}}$
$+\underbrace{\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(\bar{p}_{|\mathcal{D}_{N}}^{*}(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x}))\big]}_{\text{data uncertainty}}.$ (4)
@@
Finally, the second term in Eqn. 2.2 reflects the generalised bias, which is the difference between the mean predicted distribution $\bar{p}^{*}(y|\boldsymbol{x})$ under data and model uncertainty and the true conditional distribution $p(y|\boldsymbol{x})$. For universal approximators, such as many types of deep neural networks, the model uncertainty is zero [32, 33, 34]. Such models are asymptotically unbiased, but in the finite data regime, a bias is unavoidable to make learning possible [20]. For more restrictive hypothesis spaces, the bias even does not disappear when the training dataset becomes infinitely large. Thus, in both cases, the bias term cannot be ignored when analysing epistemic uncertainty. We see that any proper scoring rule whose associated divergence can be related to a Bregman divergence allows for the split of its expected risk into an aleatoric and an epistemic part, where the epistemic part can be further split into a part representing bias, procedural and data uncertainty, respectively. In the next sections we will look at this decomposition more closely for the regression case, using mean squared error and negative log likelihood as loss functions.
@@ 3 Evaluating epistemic uncertainty via statistical simulations
3.1 The reference distribution
Quantifying epistemic uncertainty objectively is challenging, and evaluating the correctness of such quantifications is equally difficult. While calibration tests exist for aleatoric uncertainty in probabilistic models [35], epistemic uncertainty lacks established evaluation protocols and is often assessed indirectly through downstream tasks like out-of-distribution detection [5, 22, 36], robustness to adversarial attacks [37, 38], and active learning [39]. Correctness of second-order epistemic uncertainty representations can only be evaluated through statistical simulations, in which bias and variance can be analysed individually. To this end, Jürgens et al. [24] recently introduced the notion of a reference distribution, which is the theoretical distribution over first-order models if training data could be resampled. Here we further extend this notion by taking the variance due to procedural uncertainty into account:
Definition 1.
@@ 3.1 The reference distribution
$q_{N,\gamma}(\boldsymbol{\theta}|\boldsymbol{x},\phi):=\mathbb{P}(\boldsymbol{\hat{\theta}}_{N,\gamma}(x)=\boldsymbol{\theta})=\int_{\Gamma}\int_{\mathcal{X}\times\mathcal{Y}}\mathbb{I}\Big[(\boldsymbol{\hat{\theta}}_{N,\gamma}(x)=\boldsymbol{\theta})\Big]dP^{N}dP^{\gamma},$ (5)
@@
Intuitively, the reference distribution describes the uncertainty in the first-order parameters, induced by having a limited dataset and procedural uncertainty due to different weight initializations. In simulation studies, one can estimate the reference distribution by drawing $n_{d}$ training datasets $\mathcal{D}_{N_{1}},\dots,\mathcal{D}_{N_{n_{d}}}$ and $n_{\gamma}$ procedural parameters $\gamma_{1},\dots,\gamma_{n_{\gamma}}$ from $P^{N}$ and $P^{\gamma}$, respectively, and calculating the empirical distribution function. The (generalized) variance of the reference distribution describes the variability due to limited data expressivity as well as the procedural variance, and can be decomposed by means of the decomposition in Eqn. 2.2. In addition, in statistical simulations, where the data-generating process $p(y|\boldsymbol{x})$ is known, the bias of the reference distribution can also be measured.
3.2 Bias-variance decomposition in the regression case
We start this section by deriving expressions for aleatoric and epistemic uncertainty estimates according to Section 2.2 for the regression case, which is the setting we focus on for the remainder of the paper. Regression has the advantage over classification that our key findings can be more easily and intuitively visualized, and has also not been studied as extensively as the classification case for uncertainty disentanglement.
We consider the following data-generating process: $y_{i}=f(\boldsymbol{x}_{i})+\epsilon_{i}$, where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(\boldsymbol{x}_{i}))$ is data-dependent (heteroscedastic) noise, and $f$ is a deterministic function, such that $p(\cdot|\boldsymbol{x})$ forms the density of a Gaussian distribution with mean $f(\boldsymbol{x})$ and variance $\sigma^{2}(\boldsymbol{x})$. For the case where the loss function is the mean squared error loss, $\ell_{\text{MSE}}(\hat{y},y)=(\hat{y}-y)^{2}$, Eqn. 2.2 then becomes the classical bias-variance decomposition:
$\mathbb{E}_{y|\boldsymbol{x}}\big[\mathbb{E}_{\mathcal{D}_{N},\gamma}(\hat{y}-y)^{2}|\boldsymbol{x}\big]=\sigma^{2}(\boldsymbol{x})+\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})+\text{bias}^{2}\big[\hat{y}|\boldsymbol{x}\big],$ (6)
where the variance of the prediction $\hat{y}$ can be further split into procedural and data uncertainty by means of the law of total variance:
$\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})=\mathbb{E}_{\mathcal{D}_{N}}\big[\text{Var}_{\gamma}(\hat{y}|\boldsymbol{x})|\mathcal{D}_{N}\big]+\text{Var}_{\mathcal{D}_{N}}(\mathbb{E}_{\gamma}[\hat{y}|\boldsymbol{x}]|\mathcal{D}_{N}).$ (7)
3.3 Simulating data and procedural uncertainty
@@
Figure 2: Estimated procedural and data uncertainty of the reference distribution for increasing sample sizes. The reference distribution is estimated by resampling training data and procedural parameters, using $n_{\gamma}=10$ and $n_{d}=20$. Procedural and data uncertainty are calculated using the decomposition formula in Eqn. 2.2. Even when the model predictions (complete line) have an important deviation (bias) from the ground-truth expected value (dashed line), the epistemic uncertainty estimate decreases with more data.
In the following we illustrate how one can use the reference distribution in Def. 1 to calculate expressions for data- and procedural uncertainty in the regression case. To this end, we adapt the synthetic regression setting from Rossellini et al. [40]. Precisely, we generate datasets $\mathcal{D}_{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where
$x_{i}\sim Beta(1.2,0.5),\quad y_{i}=\sin(\frac{1}{5(x_{i}+0.16)^{3}})+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(x_{i})),$ (8)
with $\sigma^{2}(x)=x^{4}$. The setup can be seen in Figure 1(a). This dataset highlights the differences between aleatoric and epistemic uncertainty: for low $x$ values, the complexity of the underlying data-generating function is high, while data is limited, leading to high epistemic uncertainty. On the other hand, for $x$ close to $1$, aleatoric uncertainty is high, while simultaneously having more data points and a less complex underlying function, implying low epistemic uncertainty. In order to estimate the reference distribution we train models with $n_{d}$ training datasets and $n_{\gamma}$ random weight initializations. As the underlying model class for obtaining the individual risk minimizers $\hat{p}_{\gamma,N}$, we use an MLP. Details about the experimental setup are discussed in Appendix A.1111The implementation for all our examples can be found in this GitHub repository.. Individual estimates for data- and procedural uncertainty can be obtained by the decomposition formula in Eqn. 2.2. The bias can be calculated as the absolute distance between the mean prediction and the underlying ground truth conditional expectation.
@@
Figure 2 shows the obtained procedural and data uncertainty of the reference distribution for different training dataset sizes. First, we see that the total variance of the reference distribution decreases when sample size increases, supporting the claim that the reference distribution properly represents data uncertainty. Second, we see that the estimate of the procedural uncertainty also decreases with an increasing sample size, which is not unexpected, because initial procedural configurations become less important when more data is observed. Third, in the region with high epistemic uncertainty, the reference distribution has a very large bias, due to the underlying complexity of the data generating function. The bias is in fact so high that aleatoric uncertainty cannot be correctly estimated anymore. This last finding confirms our central position: Any claim to properly represent epistemic uncertainty must examine not only how different sources of uncertainty are modeled, but also whether bias is systematically distorting the partitioning of uncertainty. When the bias is high, estimates of both epistemic and aleatoric uncertainty can be completely misleading.
@@ 4 Which part of epistemic uncertainty do current methods capture?
Throughout this section we evaluate methods for the heteroscedastic regression setting using the classical variance-based split of the predictive variance as in Eqn. 7. We hereby focus on the mean squared error case of the more general pointwise decomposition in Section 2.2, which is the most commonly used decomposition for regression in practice. Departing from a Bayesian perspective, we first describe in this section how various second-order methods claim to obtain disentangled estimates for aleatoric and epistemic uncertainty. Secondly, we analyse two representative methods that approximate the posterior distribution over the parameters in more detail, illustrating that they tackle (and neglect) different sources of epistemic uncertainty. In the last part of this section as well as in Appendix A.1, we present experimental results for eight selected methods, highlighting that all of them have difficulties to model the different sources of epistemic uncertainty in a complete manner.
@@ 4.1 Second-order uncertainty quantification methods
Bayesian methods form a principled approach to quantify and interpret epistemic uncertainty by modelling a prior and posterior distribution over the model parameters and computing the second-order distribution via Eqn. 1. For the regression setup that we consider in this section, a common approach is to use Heteroscedastic Gaussian Processes (GP, [19, 41]), which place priors on the parameters $f(\boldsymbol{x})$ and $\sigma^{2}(\boldsymbol{x})$, thereby analytically computing the predictive distribution using two Gaussian processes. For Bayesian neural networks, which have gained popularity in recent years, the posterior and second-order distribution cannot be analytically computed, due to the intractability of $p(\boldsymbol{\phi}|\mathcal{D}_{N})\propto p(\mathcal{D}_{N}|\boldsymbol{\phi})p(\boldsymbol{\phi})$ induced by the hypothesis space [1]. Due to the impossibility of performing exact inference in Bayesian neural networks, one typically obtains $d$ different samples $\{\boldsymbol{\phi}_{j}\}_{j=1}^{d}$ from $p(\boldsymbol{\phi}|\mathcal{D}_{N})$ by use of approximation techniques that aim to find a posterior distribution via sampling methods. Here, the second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ of Eqn. 1 is then approximated by Monte Carlo integration:
@@
In the experiments below, we employ three of the most common methods to obtain $\{\boldsymbol{\phi}_{i}\}_{j=1}^{M}$. Variational inference, (VI, [42, 43]) considers a variational (simpler) distribution $q(\boldsymbol{\phi})$ and optimizes the lower bound of the marginal likelihood (ELBO). Hamiltonian Monte Carlo (HMC) uses Hamiltonian dynamics to sample $\{\boldsymbol{\phi}_{j}\}$ from the true posterior. Laplace approximation (LA, [44]) builds a Taylor expansion around the maximum-a-posteriori estimate $\boldsymbol{\phi}_{\text{MAP}}$, yielding a Gaussian approximation. Ensemble methods are also often interpreted as performing approximate Bayesian inference by finding multiple modes of $p(\boldsymbol{\phi}|\mathcal{D}_{N})$ after training multiple independent models. [45, 46]. In this work, we consider three types of ensembles. Deep Ensembles (DE, [9]), train $d$ networks on the same dataset $\mathcal{D}_{N}$, but with different random initialization of the weights $\boldsymbol{\phi}$. Similarly, in Bootstrap Ensembles (BE), $M$ networks are trained using different weight initializations, but each network uses a bootstrap sample $\mathcal{D}^{(j)}\subseteq\mathcal{D}_{N}$. While Bootstrap Ensembles have been shown to perform worse than training $d$ networks using Deep Ensembles [9], we include them in our analysis to estimate the effect of data uncertainty on the epistemic uncertainty estimates. Monte Carlo Dropout (MCD, [47]) can also be understood as a Bayesian approximation, where one trains a single neural network with dropout rate $\rho$ and performs multiple stochastic forward passes at inference time with the same dropout rate. In a different direction, evidential deep learning methods learn directly the second-order distribution over the predictive parameters $\theta$. Deep evidential regression (DER, [12]) learns hyperparameters of a conjugate prior (e.g. Normal-Inverse-Gamma for the univariate normal distribution) directly using a single network trained by second-order loss minimization.
4.2 Uncertainty disentanglement
From a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$, one can decompose its representation of uncertainty into an estimate of aleatoric and epistemic uncertainty. Proposed measures focus on entropy [3], distance to reference sets [10], or proper scoring rules [48, 49]. The most commonly used one for regression is arguably the variance-based decomposition [3, 10], which considers the variance of the prediction $\hat{y}$, obtained via model averaging as in Eqn. 1. The latter can be decomposed into (estimates of) an aleatoric and an epistemic part:
@@ 4.3 Two representative cases
Example 1 (Deep Ensembles purely capture procedural uncertainty.).
Let $\gamma\sim P^{\gamma}$ denote the random seed (weight initialization, data ordering etc.) that differentiates each ensemble member’s training, which in turn are trained on the same dataset $\mathcal{D}_{N}$. For a Deep Ensemble of size $d$, each ensemble member’s predictive parameters are indexed over the set of random seeds $\{\gamma_{1},\dots,\gamma_{d}\}$, obtaining a set of predictions $\{\boldsymbol{\theta}_{\gamma_{i}}(\boldsymbol{x})=(f_{\gamma_{i}}(\boldsymbol{x}),\sigma_{\gamma_{i}}^{2}(\boldsymbol{x}))\}_{i=1}^{d}$. The empirical second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ can be approximated as a mixture of equally weighted Dirac measures
@@
where $\hat{y}_{N,\gamma}(\boldsymbol{x}):=f_{\gamma}(\boldsymbol{x})$ corresponds to the prediction of an individual ensemble member trained with random initialization $\gamma$ on $\mathcal{D}_{N}$ and $\hat{y}_{N}(\boldsymbol{x}):=\hat{f}(\boldsymbol{x})$ of the aggregated ensembles. Hence, we see that the predicted epistemic uncertainty corresponds to the procedural uncertainty as illustrated in Figure 1(b). Model uncertainty through bias as well as data uncertainty are not represented.
@@
Example 2 (DER may fail to provide faithful epistemic uncertainty estimates.).